# FedReFT: Federated Representation Fine-Tuning with All-But-Me Aggregation

### v1

@@ Abstract
[…]
@@
Fatema Siddika1thanks: Equal contribution, Md Anwar Hossen [^1] footnotemark: 1, J. Pablo Muñoz2, Tanya Roosta3, Anuj Sharma1, Ali Jannesari1 1Iowa State University 2Intel Labs, 3University of California, Berkeley {fatemask, manwar, anujs, jannesar}@iastate.edu jpablomch@gmail.com, tanya.roosta@gmail.com
@@ 1 Introduction
Fine-tuning has emerged as a core strategy for adapting large language models (LLMs) to various downstream tasks, allowing for a broad generalization from minimal task-specific data Ding et al. (2023); Ziegler et al. (2019). However, traditional fine-tuning is computationally expensive and memory-intensive, which poses scalability challenges. This is further amplified in resource-constrained environments, such as smartphones, where full model updates are often infeasible due to limited resources. To address these challenges, parameter-efficient fine-tuning (PEFT) methods such as Adapter Tuning Houlsby et al. (2019), BitFit Zaken et al. (2022), Prefix Tuning Li and Liang (2021), Prompt Tuning Lester et al. (2021), and Low-Rank Adaptation (LoRA) Hu et al. (2021a), have been proposed, significantly reducing the cost of adaptation by updating only a small subset of model weights.
@@
While PEFT typically modifies model weights, recent interpretability research highlights the potential of hidden representations, which encode rich semantic information. Representation Fine-Tuning (ReFT) Wu et al. (2024b) leverages this by directly intervening in hidden layers, achieving stronger performance than methods like LoRA. Despite ReFT’s success in centralized settings, it has yet to be adapted for FL, where challenges such as data heterogeneity, varying model capacities, and limited computational resources complicate aggregation and reduce effectiveness. In order to study the challenges of representation-level fine-tuning under heterogeneous federated settings, and to evaluate the effectiveness of our proposed aggregation strategy, we put forward the following research questions:
@@
1) How can we aggregate representation-level updates in Federated Learning without compromising semantic alignment across task-heterogeneous clients?
@@
2) Is weighted averaging sufficient for aligning semantically rich hidden representations, or is a more robust and personalized strategy needed to preserve local semantics while leveraging global knowledge?
@@
To address these questions, we introduce Federated Representation Fine-Tuning (FedReFT), a novel framework for personalized and parameter-efficient federated representation fine-tuning. FedReFT builds on the core idea of ReFT by enabling clients to inject lightweight intervention components (i.e., sparse low-rank matrices $W$, $R$, $b$) directly into hidden representations, making it particularly useful for edge devices with limited resources. To mitigate the degradation in semantic alignment caused by naive aggregation such as vanilla weighted average aggregation (FedAvg) McMahan et al. (2017), we propose the All-But-Me (ABM) aggregation strategy. Instead of averaging all client updates uniformly, ABM constructs a personalized global intervention for each client by computing the geometric median over updates from all other clients.
The key contributions of our work are as follows: Contribution 1: We address a critical gap in the utilization of ReFT in FL setting by introducing a novel aggregation strategy, All-But-Me (ABM), specifically designed for low-rank, representation-level interventions. ABM addresses the challenge of semantic misalignment caused by naive aggregation strategies, preserving client-specific semantics while enabling stable collaboration.
@@
Contribution 2: We propose FedReFT, a novel framework for personalized and parameter-efficient federated fine-tuning based on representation-level interventions. Contribution 3: We evaluate the framework by simulating task heterogeneity, i.e., assigning different tasks to clients, all derived from a common dataset. This setup mimics real-world scenarios where clients pursue distinct objectives over structurally similar data, allowing us to evaluate the effectiveness of FedReFT and ABM under realistic conditions. Paper outline: The remainder of the paper is organized as follows. Section 2 formally defines the problem setting, the motivation behind our work and the challenges posed by heterogeneous FL setting. Section 3 details our methodology, including the FedReFT mechanism and the All-But-Me (ABM) aggregation strategy. Section 4 presents our experimental results and evaluates the effectiveness of our approach on multiple benchmarks. Section 5 concludes the paper with key insights and future directions. We defer additional details to the appendices.
@@
Figure 1: Illustration of the relationship between the average accuracy and trainable parameter (in $\%$)for various federated PEFT methods on Commonsense, Arithmetic, and GLUE benchmarks using LLaMA-3.2B, LLaMA-3 8B, and RoBERTa-large models, respectively. FedReFT achieves competitive or state-of-the-art performance while training significantly fewer parameters, resulting in improved communication efficiency and reduced transmission cost in federated learning settings.
@@ 2 Problem Formulation and Motivation
In this section, we present the motivation to apply ReFT in FL and formalize the problem of adapting personalized representation. Although ReFT offers parameter-efficient updates in the representation space, its application in FL faces key challenges, including task heterogeneity, semantic misalignment, and unstable aggregation among heterogeneous clients. We highlight these challenges and formulate the objective of enabling parameter-efficient and semantically aligned adaptation in FL using ReFT. Challenge 1: LoReFT in FL Settings: ReFT  Wu et al. (2024b) offers an attractive alternative by modifying hidden activations instead of model weights. By intervening directly in structured semantic subspaces, ReFT supports interpretable, modular, and task-aligned adaptation, particularly advantageous in task-heterogeneous FL settings. However, full ReFT incurs considerable communication costs and poses integration challenges when clients use different model capacities or architectures.
To bridge this gap, we adopt Low-Rank Linear Subspace ReFT (LoReFT)Wu et al. (2024b), a lightweight ReFT variant that constrains interventions to a learnable low-rank subspace. This design significantly reduces overhead while maintaining semantic control, making it a promising candidate for FL. We follow the LoReFT intervention formulation from Wu et al. (2024b) on hidden representations $h\in\mathbb{R}^{d}$ which is defined as:
@@
where, $W\in\mathbb{R}^{r\times d}$ is a low-rank projection matrix with $d$ as the representation dimension and $r$ as the subspace intervention dimension, $R\in\mathbb{R}^{r\times d}$ is a low-rank projection matrix with orthonormal rows, and $b\in\mathbb{R}^{r}$, with $r\ll d$. This structure, inspired by Distributed Interchange Intervention (DII) Geiger et al. (2024), enables semantically grounded, low-rank adaptation suitable for scalable and privacy-preserving FL. Despite its efficiency, applying LoReFT in FL raises several non-trivial challenges: LoReFT modifies internal representations that are sensitive to client-specific data distributions. Aggregating these interventions naïvely using FedAvg can cause semantic interference or collapse. Without global synchronization, low-rank updates may evolve in divergent directions, especially when tasks are dissimilar. Applying shared LoReFT interventions across clients risks overfitting to shared patterns while ignoring local semantics. Considering all these challenges, the major research question is:
@@
FedReFT uses All-But-Me(ABM) aggregation to robustly combine intervention parameters while preserving personalization in heterogeneous FL. Challenge 2: Federated Fine-Tuning under Task Heterogeneity: A central motivation of our work is to address task heterogeneity in real-world FL, where clients perform fundamentally different tasks rather than optimizing a shared objective. For example, clients may work on distinct reasoning tasks within natural language QA that demand different semantic skills. While centralized fine-tuning has proven effective for such tasks, it assumes access to all data, which is unrealistic in decentralized settings. In FL, each client sees only a local, task-specific subset of the broader reasoning space, leading to highly heterogeneous training distributions, a common challenge in multi-department or cross-domain deployments. This raises the question:
How can we learn a global representation that generalizes across tasks when each client trains only on a fragment of the broader task distribution?
Standard methods like FedAvg McMahan et al. (2017) struggle in this regime, as they average semantically misaligned updates, often resulting in degraded performance or collapsed representations. Formally, let each client $i$ have a dataset $\mathcal{T}_{i}=\{X_{i},Y_{i}\}$ and optimize a personalized model $\boldsymbol{\theta}_{i}$ by solving:
@@
where $\mathcal{L}$ is the task-specific loss and $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{i}\}_{i=1}^{N}$ is the set of client-specific models. Our proposed method, FedReFT, can successfully address this research challenge. FedReFT enables scalable, personalized representation learning across heterogeneous tasks, allowing global reasoning capabilities to emerge from decentralized, task-specific updates.
Table 1: Comparison of different aggregation methods of GLUE task on ROBERTa, Commonsense and Arithmetic reasoning task on LLaMA-2 7B.
Task FedAvg (%) FedReFT (%)
Commonsense 70.16 70.77
Arithmetic 15.36 17.21
GLUE 88.17 89.77
Challenge 3: Learnable Parameter Sharing with the Server: When applying ReFT Wu et al. (2024b) in a FL setting from the perspective of learnable parameter sharing, a fundamental question is:
@@
Figure 2: FedReFT with ABM Aggregation. Clients cross-task demonstrate personalization while maintaining alignment with the global representation. (1)-(2): Each client applies LoReFTWu et al. (2024b) interventions to train learnable parameter {$W$, $R$, bb} to modify hidden representations $h$ in a low-rank edit subspace. (3): Clients fine-tune {$W$, $R$, bb} locally and partially fuse received All-But-Me aggregated updates with their own. (4): The server performs ABM aggregation using the geometric median over other clients’ intervention parameters to generate $W_{k}^{\text{ABM}},R_{k}^{\text{ABM}},b_{k}^{\text{ABM}}$.
@@ 3 Methodology
In this section, we introduce FedReF+, designed to address the challenges we discussed in the previous section. An illustrative overview of FedReFT is shown in Figure 2.
@@ 3.1 Intervention Parameter Sharing Strategies
To reduce communication overhead while maintaining personalization, we explore three strategies for sharing local intervention parameters with the server. These strategies are summarized in Table 2 and represent different trade-offs between expressiveness and communication efficiency:
1) Full Intervention Sharing: $\mathbf{W}\in\mathbb{R}^{r\times d}$, $\mathbf{R}\in\mathbb{R}^{r\times d}$, $\mathbf{b}\in\mathbb{R}^{r}$} This strategy shares the complete set of intervention parameters, capturing client-specific compression ($\mathbf{W}$), transformation ($\mathbf{R}$), and translation ($\mathbf{b}$). It enables the most accurate reconstruction of local updates and yields the best global performance, especially under high heterogeneity.
2) No Bias Sharing: $\mathbf{W}\in\mathbb{R}^{r\times d}$, $\mathbf{R}\in\mathbb{R}^{r\times d}$} This variant omits the bias term $\mathbf{b}$ but retains the directional transformation via $\mathbf{W}$ and $\mathbf{R}$. While it allows the server to align low-rank subspace transformations across clients, it lacks the ability to model per-dimension translation shifts, which can hinder fine-grained personalization. 3) No $\mathbf{W}$ Sharing: $\mathbf{R}\in\mathbb{R}^{r\times d}$, $\mathbf{b}\in\mathbb{R}^{r}$} This configuration excludes $\mathbf{W}$, giving the server access only to the reconstruction and shift parameters. Without knowledge of how the local signals were encoded, the server’s ability to interpret or align updates is severely limited. The $\{\mathbf{W},\mathbf{R},\mathbf{b}\}$ strategy provides the highest fidelity for aggregation, $\{\mathbf{W},\mathbf{R}\}$ offers a balanced compromise, and $\{\mathbf{R},\mathbf{b}\}$ prioritizes communication efficiency at the cost of semantic alignment and global performance.
Table 2: Performance vs. parameter efficiency for different LoReFT sharing strategies (Uplink) for $C$ clients on commonsense reasoning task following the second experiment design. GLUE task on ROBERTa, Arithmetic and Commonsense on LLaMa-2 7B model.
Task Strategy TP(% $\downarrow$) Score
GLUE W,R,b 0.01384 94.31
W,R 0.01383 64.03
R,b 0.00693 74.12
Arithmetic W,R,b 0.03114 29.01
W,R 0.03114 26.13
R,b 0.01557 25.77
Commonsense W,R,b 0.03114 73.82
W,R 0.03114 70.63
R,b 0.01557 68.02
@@ 3.2 Intervention Design for Federated Classification Tasks
Following the formulation in ReFT Wu et al. (2024b), for a given client, we define the classification head $H_{\psi}$ with parameters $\psi=\{W_{o},b_{o},W_{d},b_{d}\}$ operates on the CLS token representation $z\in\mathbb{R}^{d}$ from the final layer:
@@
3.3 All-But-Me (ABM) Aggregation
In heterogeneous FL, the integration of shared knowledge without compromising local task-specific adaptation remains a core challenge. Standard aggregation methods such as FedAvg McMahan et al. (2017), which averages client models into a single global model, are often suboptimal in non i.i.d. scenarios. They risk overwriting valuable client-specific representations and rely on fixed mixing weights that may further reduce personalization. Table 1 depicts the comparison. To overcome these limitations, we propose the All-But-Me (ABM) aggregation strategy. Instead of initializing clients with a global model, each client continues to update its local parameters while partially incorporating knowledge aggregated from other clients. Specifically, each client $k$ receives a robustly aggregated set of intervention parameters $\{\mathbf{W}^{\text{ABM}}_{k},\mathbf{R}^{\text{ABM}}_{k},\mathbf{B}^{\text{ABM}}_{k}\}$, calculated from the updates of all other clients using a geometric median:
$\mathbf{W}_{k}^{\text{ABM}}=\text{ABM}\left(\left\{\mathbf{W}_{m}^{\text{local}}\right\}_{m\neq k}\right),$ (5)
$\mathbf{R}_{k}^{\text{ABM}}=\text{ABM}\left(\left\{\mathbf{R}_{m}^{\text{local}}\right\}_{m\neq k}\right),$
$\mathbf{B}_{k}^{\text{ABM}}=\text{ABM}\left(\left\{\mathbf{B}_{m}^{\text{local}}\right\}_{m\neq k}\right).$
The client then performs a personalized update by interpolating between its local parameters and the ABM-aggregated ones using a mixing factor $\alpha\in[0,1]$. We have discussed in detail how the evaluation loss-based $\alpha$ tuning works in the appendix D.
$\mathbf{X}_{k}^{\text{new}}$ $=(1-\alpha)\cdot\mathbf{X}_{k}^{\text{local}}+\alpha\cdot\mathbf{X}_{k}^{\text{ABM}},$ (6)
$\mathbf{X}$ $\in\{\mathbf{W},\mathbf{R},\mathbf{B}\}.$
Before the next local training using the $\mathbf{R}_{k}^{\text{new}}$, we do the orthogonal transformation of $\mathbf{R}_{k}^{\text{new}}$ to keep the original property of $R$.
ABM via Geometric Median.
The geometric median (also known as the spatial or $L_{1}$ median) offers a robust alternative to the arithmetic mean, particularly under client heterogeneity and adversarial conditions Maronna and Martin (2006); Weiszfeld (1937). Given a set of vectors $\mathcal{S}=\{x_{1},\ldots,x_{n}\}\subset\mathbb{R}^{d}$, it is defined as:

### v2

@@ Abstract
Parameter-efficient fine-tuning (PEFT) adapts large pre-trained models by updating only a small subset of parameters. Recently, Representation Fine-Tuning (ReFT) has emerged as an effective alternative. ReFT shifts the fine-tuning paradigm from updating model weights to directly manipulating hidden representations that capture rich semantic information, and outperform state-of-the-art PEFTs in standalone settings. However, its application in Federated Learning (FL) remains challenging due to heterogeneity in clients’ data distributions, model capacities, and computational resources. To address these challenges, we introduce Federated Representation Fine-Tuning (FedReFT), a novel approach to fine-tune clients’ hidden representations. FedReFT applies sparse intervention layers to steer hidden representations directly, offering a lightweight and semantically rich fine-tuning alternative ideal for edge devices. However, representation-level updates are especially vulnerable to aggregation mismatch under different task heterogeneity, where naive averaging can corrupt semantic alignment. To mitigate this issue, we propose All-But-Me (ABM) aggregation, where each client receives the aggregated updates of others and partially incorporates them, enabling stable and personalized learning by balancing local focus with global knowledge. We further design an adaptive update strategy inspired by Test-Time Computing (TTC) to balance local and global contributions under heterogeneous conditions. FedReFT achieves state-of-the-art performance on commonsense reasoning, arithmetic reasoning, and GLUE benchmarks, while delivering $1\times$$49\times$ higher parameter efficiency compared to leading LoRA-based methods. The paper code is available at Anonymous Repository
@@
Fatema Siddika1thanks: Equal contribution, Md Anwar Hossen [^1] footnotemark: 1, Juan Pablo Munoz2, Tanya Roosta3 Anuj Sharma1, Ali Jannesari1 1Iowa State University, Ames, USA 2Maro Systems, USA 3University of California, Berkeley {fatemask, manwar, anujs, jannesar}@iastate.edu pablo.munoz@maro-systems.com, tanya.roosta@gmail.com
@@ 1 Introduction
Fine-tuning has emerged as a core strategy for adapting large language models (LLMs) to various downstream tasks, allowing for a broad generalization from minimal task-specific data ding2023parameter; ziegler2019fine. However, traditional full fine-tuning is computationally expensive and memory-intensive, which poses scalability challenges. This is further amplified in resource-constrained environments, such as smartphones, where full model updates are often infeasible due to limited resources. To address these challenges, parameter-efficient fine-tuning (PEFT) methods, such as Adapter Tuning houlsby2019parameter, BitFit zaken2022bitfit, Prefix Tuning li2021prefix, Prompt Tuning lester2021power, and Low-Rank Adaptation (LoRA) Hu21, have been proposed. These methods significantly reduce the cost of adaptation by updating only a small subset of model weights.
@@
While PEFT typically modifies model weights, recent interpretability research highlights the potential of hidden representations, which encode rich semantic information. ReFT Wu2024 leverages this by training small interventions that act as lightweight transformations on the model’s internal representations, steering model behavior for downstream tasks without altering the original weights. This representation-level adaptation enables ReFT to achieve stronger performance than weight-based methods such as LoRA. Despite ReFT’s success in centralized settings, it has yet to be adapted to FL setting, where challenges such as data heterogeneity, varying model capacities, and limited computational resources complicate aggregation and reduce effectiveness. To investigate the challenges of representation-level fine-tuning in heterogeneous federated settings and assess the effectiveness of our proposed aggregation strategy, we formulate the following research questions:
@@
RQ1: How can representation-level updates be aggregated in FL while preserving semantic alignment across task-heterogeneous clients?
@@
RQ2: Is simple weighted averaging sufficient for aligning semantically rich, hidden representations, or is a more robust and personalized strategy required to maintain local semantics while leveraging global knowledge?
@@
To address these challenges, we propose Federated Representation Fine-Tuning (FedReFT), a framework for personalized and parameter-efficient federated fine-tuning. FedReFT extends ReFT by injecting lightweight intervention components (sparse low-rank matrices $W$, $R$, $b$) directly into hidden representations, making it suitable for resource-limited edge devices. To avoid semantic misalignment from naive aggregation such as FedAvg mcmahan2017communication, we introduce All-But-Me (ABM) aggregation, which builds a global aggregated intervention by computing the geometric median over updates from all other clients. The key contributions of our work are as follows:
Contribution 1: We address a critical gap in adapting ReFT to federated learning by introducing FedReFT, the first framework that enables personalized and parameter-efficient fine-tuning through sparse representation-level interventions, making it effective for resource-constrained clients.
Contribution 2: To support this framework, we propose the All-But-Me (ABM) aggregation strategy, specifically designed for representation-level interventions. ABM mitigates semantic misalignment caused by naive averaging and enables stable, personalized collaboration under heterogeneous conditions. Furthermore, FedReFT incorporates an adaptive mixing mechanism inspired by Test-Time Computing (TTC) to dynamically learn how to combine local and ABM intervention parameters for each client.
@@
Contribution 3: We evaluate the framework by simulating task heterogeneity, i.e., assigning different tasks to clients, all derived from a common dataset. This setup mimics real-world scenarios where clients pursue distinct objectives over structurally similar data, allowing us to evaluate the effectiveness of FedReFT and ABM under realistic conditions. The rest of the paper is organized as follows. Section 2 defines the problem and challenges in heterogeneous FL. Section 3 presents our FedReFT framework and ABM aggregation. Section 4 reports experimental results, and Section 5 concludes with insights and future directions. Additional details are provided in the appendix.
@@
Figure 1: Average accuracy vs. trainable parameters (%) for federated PEFT methods on Arithmetic, Commonsense, and GLUE benchmarks using LLaMA-3 8B, LLaMA-3.2B, and RoBERTa-large models, respectively. FedReFT attains state-of-the-art accuracy while training far fewer parameters, improving communication efficiency and reducing transmission cost in FL.
@@ 2 Problem Formulation and Motivation
In this section, we motivate applying ReFT to FL and formalize the problem of personalized representation adaptation. While ReFT enables parameter-efficient updates in the representation space, its use in FL is hindered by task heterogeneity, semantic misalignment, and unstable aggregation. Therefore, we define the objective as achieving parameter-efficient and semantically aligned adaptation in FL with ReFT.
Challenge 1: LoReFT in FL Settings (RQ1): ReFT  Wu2024 offers an attractive alternative by modifying hidden activations instead of model weights. By intervening directly in structured semantic subspaces, ReFT supports interpretable, modular, and task-aligned adaptation, particularly advantageous in task-heterogeneous FL settings. However, deploying full ReFT in federated settings can be challenging, as transmitting high-dimensional representation updates increases communication overhead. Moreover, when clients employ models with varying capacities or architectures, aligning representation spaces becomes nontrivial, making integration across clients difficult.
To bridge this gap, we adopt Low-Rank Linear Subspace ReFT (LoReFT)Wu2024, which is a lightweight ReFT variant that constrains interventions to a learnable low-rank subspace. This design significantly reduces overhead while maintaining semantic control, making it a promising candidate for FL. We follow the LoReFT intervention formulation from Wu2024 on hidden representations $h\in\mathbb{R}^{d}$ which is defined as:
@@
where, $W\in\mathbb{R}^{r\times d}$ is a low-rank projection matrix with $d$ as the representation dimension and $r$ as the subspace intervention dimension, $R\in\mathbb{R}^{r\times d}$ is a low-rank projection matrix with orthonormal rows, and $b\in\mathbb{R}^{r}$, with $r\ll d$. This structure, inspired by Distributed Interchange Intervention (DII) geiger2024finding, enables semantically grounded, low-rank adaptation suitable for scalable and privacy-preserving FL. Despite its efficiency, applying LoReFT in FL raises several non-trivial challenges: (i) LoReFT modifies internal representations that are sensitive to client-specific data distributions. Aggregating these interventions naïvely using FedAvg can cause semantic interference or collapse. (ii) Without global synchronization, low-rank updates may evolve in divergent directions, especially when tasks are dissimilar. (iii) Applying shared LoReFT interventions across clients risks overfitting to shared patterns while ignoring local semantics. Considering all these challenges, the major research question is:
@@
FedReFT uses All-But-Me(ABM) aggregation to robustly combine intervention parameters while preserving personalization in heterogeneous FL. Challenge 2: Federated Fine-Tuning under Task Heterogeneity (RQ1 & RQ2): A central motivation of our work is to address task heterogeneity in real-world FL, where clients perform fundamentally different tasks rather than optimizing a shared objective. For example, clients may work on distinct reasoning tasks within natural language question answering that demand different semantic skills. While centralized fine-tuning has proven effective for such tasks, it assumes access to all data, which is unrealistic in decentralized settings. In FL, each client sees only a local, task-specific subset of the broader reasoning space, leading to highly heterogeneous training distributions, a common challenge in multi-department or cross-domain deployments. This raises the question: How can we learn a global representation that generalizes across tasks when each client trains only on a fragment of the broader task distribution? Standard methods like FedAvg mcmahan2017communication struggle in this regime, as they average semantically misaligned updates, often resulting in degraded performance or collapsed representations. Formally, let each client $i$ have a dataset $\mathcal{T}_{i}=\{X_{i},Y_{i}\}$ and optimize a personalized model $\boldsymbol{\theta}_{i}$ by solving:
@@
where $\mathcal{L}$ is the task-specific loss and $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{i}\}_{i=1}^{N}$ is the set of client-specific models. Our proposed method, FedReFT, can successfully address this research challenge. FedReFT enables scalable, personalized representation learning across heterogeneous tasks, allowing global reasoning capabilities to emerge from decentralized, task-specific updates. Challenge 3: Learnable Parameter Sharing with the Server (RQ2): When applying ReFT Wu2024 in a FL setting from the perspective of learnable parameter sharing, a fundamental question is:
@@
Figure 2: FedReFT with ABM Aggregation. Clients cross-task demonstrate personalization while maintaining alignment with the global representation. (1)-(2): Each client applies LoReFTWu2024 interventions to train learnable parameter {$W$, $R$, bb} to modify hidden representations $h$ in a low-rank edit subspace. (3): Clients fine-tune {$W$, $R$, bb} locally and partially fuse received All-But-Me aggregated updates with their own. (4): The server performs ABM aggregation using the geometric median over other clients’ intervention parameters to generate $W_{k}^{\text{ABM}},R_{k}^{\text{ABM}},b_{k}^{\text{ABM}}$.
@@ 3 Methodology
In this section, we introduce FedReFT, designed to address the challenges we discussed in the previous section. An illustrative overview of FedReFT is shown in Figure 2.
@@ 3.1 Intervention Parameter Sharing Strategies
An intervention is a small, learnable operation inserted into a model’s hidden layers that slightly modifies its internal representations, known as activations. Instead of updating the model’s weights, it transforms these hidden features using a few parameters in Equation 1; a low-rank projection matrix $\mathit{R}\in\mathbb{R}^{r\times d}$ that captures compact feature subspaces, a transformation matrix $\mathit{W}\in\mathbb{R}^{r\times d}$ that maps these subspaces back into the model’s space, and a bias vector $\mathit{b}\in\mathbb{R}^{r}$ that shifts the representation to adjust how information flows through the network. We explore different strategies for sharing local intervention parameters with the server to reduce communication overhead while maintaining personalization. These strategies are summarized in Appendix D.1 and represent different trade-offs between expressiveness and communication efficiency. After observing results and analysis, all the experiments shared the full Intervention parameters.
@@ 3.2 Intervention Design for Federated Classification Tasks
Following the formulation in ReFT Wu2024, for a given client, we define the classification head $H_{\psi}$ with parameters $\psi=\{W_{o},b_{o},W_{d},b_{d}\}$ which operates on the CLS token representation $z\in\mathbb{R}^{d}$ from the final layer:
@@
3.3 All-But-Me (ABM) Aggregation on Server
In heterogeneous FL, the integration of shared knowledge without compromising local task-specific adaptation remains a core challenge. Standard aggregation methods, such as FedAvg mcmahan2017communication, which averages client models into a single global model, are often suboptimal in non i.i.d. scenarios. They risk overwriting valuable client-specific representations and rely on fixed mixing weights that may further reduce personalization. To overcome these limitations, we propose the All-But-Me (ABM) aggregation strategy. Instead of initializing clients with a global model, each client continues to update its local parameters while partially incorporating knowledge aggregated from other clients. Specifically, each client $k$ receives a robustly aggregated set of intervention parameters $\{\mathit{W}^{\text{ABM}}_{k},\mathit{R}^{\text{ABM}}_{k},\mathit{B}^{\text{ABM}}_{k}\}$, calculated from the updates of all other clients using a geometric median:
$X=\{\mathit{W},\mathit{R},\mathit{B}\},\quad X_{k}^{\text{ABM}}=\text{ABM}(\left\{X_{m}^{\text{c}}\right\}_{m\neq k}).$ (5)
ABM via Geometric Median. The geometric median (also known as the spatial or $L_{1}$ median) offers a robust alternative to the arithmetic mean, particularly under client heterogeneity and adversarial conditions maronna2006yohai; weiszfeld1937point. Given a set of vectors $\mathcal{S}=\{x_{1},\ldots,x_{n}\}\subset\mathbb{R}^{d}$, it is defined as: